f(x,y)= -(x2-1)2-(x2y-x-1)2

Find the critical points of f (x, y) and Using the Second Partials Test (SPT), find the relative extrema’s of f (x ,y)


1
Expert's answer
2022-04-28T10:25:44-0400
"f(x,y)= -(x^2-1)^2-(x^2y-x-1)^2"

Find the critical point(s)


"f_x=-2(x^2-1)(2x)-2(x^2y-x-1)(2xy-1)"

"f_y=-2x^2(x^2y-x-1)"

"\begin{cases}
 f_x=0 \\
 f_y=0
\end{cases}"

If "x=0," then


"-2(0-1)(0)-2(0-0-1)(0-1)=-2\not=0"

If "x^2y-x-1=0," then


"-2(x^2-1)(2x)-2(0)(2xy-1)=0"

Since "x\not=0," we take "x_1=-1, x_2=1."

"x=-1"


"(-1)^2y-(-1)-1=0=>y=0"

Point "(-1, 0)."


"x=1"

"(1)^2y-1-1=0=>y=2"

Point "(1, 2)."

"f_{xx}=-12x^2+2-12x^2y^2+12xy+4y"

"f_{xy}=f_{yx}=-8x^3y+6x^2+4x"

"f_{yy}=-2x^4"

Point "(-1, 0)"

"f_{xx}(-1,0)=-10<0"


"D=\begin{vmatrix}
 -10 & 2 \\
 2 & -2
\end{vmatrix}=16>0"

There is a relative maximum at "(-1, 0)."


Point "(1, 2)"

"f_{xx}(1,2)=-26<0""D=\begin{vmatrix}
 -26 & -6 \\
 -6 & -2
\end{vmatrix}=16>0"

There is a relative maximum at "(1, 2)."